# Monotonicity properties of weighted geometric symmetrizations

### v1

@@ Abstract.
We prove new monotonicity properties for spectral radius, essential spectral radius, operator norm, Hausdorff measure of non-compactness and numerical radius of products and sums of weighted geometric symmetrizations of positive kernel operators on $L^{2}$. To our knowledge, several proved properties are new even in the finite dimensional case.
@@
weighted Hadamard-Schur geometric mean; Hadamard-Schur product; spectral radius; essential spectral radius; operator norm; Hausdorff measure of non-compactness; numerical radius; positive kernel operators
@@ 1. Introduction
for Hadamard weighted geometric mean holds for nonnegative $n\times n$ matrices $A_{1},A_{2},\ldots,A_{m}$ and nonnegative numbers $\alpha_{1}$, $\alpha_{2}$,…, $\alpha_{m}$ such that $\sum_{j=1}^{m}\alpha_{j}\geq 1$. Here $A^{(\alpha)}=[a_{ij}^{\alpha}]$ denotes the Hadamard (Schur) power of $A$ and $A\circ B=[a_{ij}b_{ij}]$ denotes the Hadamard (Schur) product of matrices $A$ and $B$. Clearly, (1.2) generalizes (1.1), since $S(A)=A^{(\frac{1}{2})}\circ(A^{T})^{(\frac{1}{2})}$. Let us point out that inequality (1.2) can straightforwardly be deduced from an earlier result by Kingman [23] and that (1.1) is a special case of earlier results by Karlin and Ost [22, Theorem 2.1 and Remark 1] and that the case $\sum_{j=1}^{m}\alpha_{j}=1$ of (1.2) was already obtained in [22, Remark 1] (in [22] these results were applied in the context of finite stationary Markov chains). Since then inequalities and equalities on Hadamard weighted geometric means and weighted geometric symmetrizations received a lot of attention and have been applied in a variety of contexts (see e.g. [15, 10, 26, 21, 12, 30, 39, 42, 13, 3, 17, 18, 31, 37, 38, 28, 32, 14, 8, 33, 34, 43, 11, 35, 36, 6, 7, 5, 24]).
@@
In [10], Drnovšek proved that in the case when $\sum_{j=1}^{m}\alpha_{j}=1$, inequality (1.2) holds also for positive compact operators on Banach function spaces. In [12], Drnovšek and the second author of the current article proved that the compactness assumption can be removed and that analogous results hold also for operator norm and also for numerical radius on $L^{2}$. Further they proved additional results for products of Hadamard weighted geometric means (see Theorem 2.1 below). In [30], the second author also showed that analogous results also hold for Hausdorff measure of non-compactness and for essential spectral radius on suitable Banach functions spaces (including $L^{2}$, see Theorem 2.1 below). In [12] and [30], also generalizations of inequality (1.1) for products and sums of geometric symmetrizations of positive kernel operators were proved (see inequalities (3.1) and (3.2) below).
@@
In [39], Shen and Huang studied weighted geometric symmetrizations $S_{\alpha}(A)=[a_{ij}^{\alpha}a_{ji}^{1-\alpha}]$ for $\alpha\in[0,1]$ and for nonnegative $n\times n$ matrices. They showed that for a given square nonnegative matrix $A$ the function $\alpha\mapsto r(S_{\alpha}(A))$ is decreasing on $[0,\frac{1}{2}]$ and increasing on $[\frac{1}{2},1]$ ([39, Theorem 3.3]). They also proved an analogous result for the operator (largest singular value) norm ([39, Theorem 2.3]). In [6, Theorem 2.7], we obtained an analogous result for the spectral radius, essential spectral radius, operator norm, Hausdorff measure of non-compactness and numerical radius of weighted geometric symmetrizations of a given positive kernel operator on $L^{2}$. In the current article we further extend a technique of Shen and Huang to obtain additional results. For instance, as a special case of our results (see Corollary 3.3 below) we show that also the function $\alpha\mapsto r(S_{\alpha}(A_{1})S_{\alpha}(A_{2}))$ is decreasing on $[0,\frac{1}{2}]$ and increasing on $[\frac{1}{2},1]$, where $A_{1}$ and $A_{2}$ are positive kernel operators on $L^{2}$ (and that the analogue of this result holds also for the essential spectral radius).
@@
The rest of the article is organized in the following way. In Section 2 we recall some definitions and results that will be needed in our proofs. In Section 3 we prove new monotonicity properties for spectral radius, essential spectral radius, operator norm, Hausdorff measure of non-compactness and numerical radius of products and sums of weighted geometric symmetrizations of positive kernel operators on $L^{2}$. The main results of this article are Theorems 3.1 and 3.5.
@@ 2. Preliminaries
If $\{f_{n}\}_{n\in\mathbb{N}}\subset M(X,\mu)$ is a decreasing sequence and $f=\inf\{f_{n}\in M(X,\mu):n\in\mathbb{N}\}$, then we write $f_{n}\downarrow f$. A Banach function space $L$ has an order continuous norm, if $0\leq f_{n}\downarrow 0$ implies $\|f_{n}\|_{L}\to 0$ as $n\to\infty$. It is well known that spaces $L^{p}(X,\mu)$, $1\leq p<\infty$, have order continuous norm. Moreover, the norm of any reflexive Banach function space is order continuous. In particular, we are interested in Banach function spaces $L$ such that $L$ and its Banach dual space $L^{*}$ have order continuous norms. Examples of such spaces are $L^{p}(X,\mu)$, $1<p<\infty$, while the space $L=c_{0}$ is an example of a non-reflexive Banach sequence space, such that $L$ and $L^{*}=l^{1}$ have order continuous norms.
By an operator on a Banach function space $L$ we always mean a linear operator on $L$. An operator $K$ on $L$ is said to be positive if it maps nonnegative functions to nonnegative ones, i.e., $KL_{+}\subset L_{+}$, where $L_{+}$ denotes the positive cone $L_{+}=\{f\in L:f\geq 0\;\mathrm{a.e.}\}$. Given operators $K$ and $H$ on $L$, we write $K\geq H$ if the operator $H-K$ is positive.
@@
where $D_{L}=\{f\in L:\|f\|_{L}\leq 1\}$. Then $\gamma(K)\leq\|K\|$, $\gamma(K+H)\leq\gamma(K)+\gamma(H)$, $\gamma(KH)\leq\gamma(K)\gamma(H)$ and $\gamma(\alpha K)=\alpha\gamma(K)$ for $\alpha\geq 0$. Also $0\leq K\leq H$ implies $\gamma(K)\leq\gamma(H)$ (see e.g. [25, Corollary 4.3.7 and Corollary 3.7.3]). Let $r_{ess}(K)$ denote the essential spectral radius of $K$, i.e., the spectral radius of the Calkin image of $K$ in the Calkin algebra. Then
@@
and $r_{ess}(K)\leq\gamma(K)$. Recall that if $L=L^{2}(X,\mu)$, then $\gamma(K^{*})=\gamma(K)$ and $r_{ess}(K^{*})=r_{ess}(K)$, where $K^{*}$ denotes the adjoint of $K$ (see e.g. [25, Proposition 4.3.3, Theorems 4.3.6 and 4.3.13 and Corollary 3.7.3], [29, Theorem 1], [24]). Note that equalities (2.2) and $r_{ess}(K^{*})=r_{ess}(K)$ are valid for any bounded operator $K$ on a given complex Banach space $L$ (see e.g. [25, Theorem 4.3.13 and Proposition 4.3.11], [29, Theorem 1]).
It is well-known that kernel operators play a very important, often even central, role in a variety of applications from differential and integro-differential equations, problems from physics (in particular from thermodynamics), engineering, statistical and economic models, etc (see e.g. [19, 35] and the references cited there). For the theory of Banach function spaces and more general Banach lattices we refer the reader to the books [41, 4, 1, 2, 25].
@@
Let us recall the following result, which was proved in [12, Theorem 2.2] and [30, Theorem 5.1 and Example 3.7] (see also e.g. [33, Theorem 2.1]).
@@
$f_{1}^{\alpha}g_{1}^{\beta}+\cdots+f_{m}^{\alpha}g_{m}^{\beta}\leq(f_{1}+\cdots+f_{m})^{\alpha}(g_{1}+\cdots+g_{m})^{\beta}$ (2.13)
@@
$(f_{11}^{\alpha_{1}}\cdots f_{1m}^{\alpha_{m}})+\cdots+(f_{n1}^{\alpha_{1}}\cdots f_{nm}^{\alpha_{m}})\leq(f_{11}+\cdots+f_{n1})^{\alpha_{1}}\cdots(f_{1m}+\cdots+f_{nm})^{\alpha_{m}}$ (2.14)
3. New results
@@
$\rho(S_{\alpha}(K_{1})+\ldots+S_{\alpha}(K_{n}))\leq\rho(K_{1}+\cdots+K_{n})$ (3.2)
The following result generalizes [6, Theorem 2.7] by extending a technique of Shen and Huang [39].
@@
Let $K_{1},\ldots K_{n}$ be positive kernel operators on $L=L^{2}(X,\mu)$. For $\rho\in\{r,r_{ess},\gamma,\|\cdot\|,w\}$ define $\rho_{n}:[0,1]\to[0,\infty)$ by

### v2

# Monotonicity properties for joint and generalized spectral radius and their essential versions of weighted geometric symmetrizations

@@ Abstract.
We prove new monotonicity properties for joint and generalized spectral radius and their essential versions of weighted geometric symmetrizations of bounded sets of positive kernel operators on $L^{2}$. To our knowledge, several proved properties are new even in the finite dimensional case.
@@
weighted Hadamard-Schur geometric mean; Hadamard-Schur product; weighted geometric symmetrizations; bounded sets of operators; positive kernel operators; joint spectral radius; generalized spectral radius; essential joint and essential generalized spectral radius; operator norm; Hausdorff measure of non-compactness; numerical radius
@@ 1. Introduction
for Hadamard weighted geometric mean holds for nonnegative $n\times n$ matrices $A_{1},A_{2},\ldots,A_{m}$ and nonnegative numbers $\alpha_{1}$, $\alpha_{2}$,…, $\alpha_{m}$ such that $\sum_{j=1}^{m}\alpha_{j}\geq 1$. Here $A^{(\alpha)}=[a_{ij}^{\alpha}]$ denotes the Hadamard (Schur) power of $A$ and $A\circ B=[a_{ij}b_{ij}]$ denotes the Hadamard (Schur) product of matrices $A$ and $B$. Clearly, (1.2) generalizes (1.1), since $S(A)=A^{(\frac{1}{2})}\circ(A^{T})^{(\frac{1}{2})}$. Let us point out that inequality (1.2) can straightforwardly be deduced from an earlier result by Kingman [21] and that (1.1) is a special case of earlier results by Karlin and Ost [20, Theorem 2.1 and Remark 1] and that the case $\sum_{j=1}^{m}\alpha_{j}=1$ of (1.2) was already obtained in [20, Remark 1] (in [20] these results were applied in the context of finite stationary Markov chains). Since then inequalities and equalities on Hadamard weighted geometric means and weighted geometric symmetrizations received a lot of attention and have been applied in a variety of contexts (see e.g. [15, 11, 13, 28, 35, 3, 33, 34, 29, 12, 31, 32, 6, 7, 5, 22, 8] and also a more detailed list of references in [8].
@@
In [11], Drnovšek proved that in the case when $\sum_{j=1}^{m}\alpha_{j}=1$, inequality (1.2) holds also for positive compact operators on Banach function spaces. In [13], Drnovšek and the second author of the current article established that the compactness assumption can be removed and that analogous results hold also for operator norm and also for numerical radius on $L^{2}$. Further they proved additional results for products of Hadamard weighted geometric means (see Theorem 2.1 below). In [28], the second author also showed that analogous results also hold for Hausdorff measure of non-compactness and for essential spectral radius on suitable Banach functions spaces (including $L^{2}$, see Theorem 2.1 below). In [13] and [28], also generalizations of inequality (1.1) for products and sums of geometric symmetrizations of positive kernel operators were proved (see inequalities (2.24) and (2.25) below).
@@
In [35], Shen and Huang studied weighted geometric symmetrizations $S_{\alpha}(A)=[a_{ij}^{\alpha}a_{ji}^{1-\alpha}]$ for $\alpha\in[0,1]$ and for nonnegative $n\times n$ matrices. They showed that for a given square nonnegative matrix $A$ the function $\alpha\mapsto r(S_{\alpha}(A))$ is decreasing on $[0,\frac{1}{2}]$ and increasing on $[\frac{1}{2},1]$ ([35, Theorem 3.3]). They also proved an analogous result for the operator (largest singular value) norm ([35, Theorem 2.3]). In [6, Theorem 2.7], we obtained an analogous result for the spectral radius, essential spectral radius, operator norm, Hausdorff measure of non-compactness and numerical radius of weighted geometric symmetrizations of a given positive kernel operator on $L^{2}$. In [8] we further extended a technique of Shen and Huang to obtain additional results. For instance, it was proved in [8, Corollary 3.3] that also the function $\alpha\mapsto r(S_{\alpha}(A_{1})S_{\alpha}(A_{2}))$ is decreasing on $[0,\frac{1}{2}]$ and increasing on $[\frac{1}{2},1]$, where $A_{1}$ and $A_{2}$ are positive kernel operators on $L^{2}$ (and that the analogue of this result holds also for the essential spectral radius). In the current article we generalize the results of [8] to the setting of bounded sets of positive kernel operators on $L^{2}$ by establishing new results for joint and generalized spectral radius, for essential versions of joint and generalized spectral radius and for operator norm, Hausdorff measure of non-compactness and numerical radius of such sets.
@@
The rest of the article is organized in the following way. In Section 2 we recall some definitions and results that will be needed in our proofs and in Section 3 we obtain new results. The main results of this article are Theorems 3.4 and 3.10.
@@ 2. Preliminaries
If $\{f_{n}\}_{n\in\mathbb{N}}\subset M(X,\mu)$ is a decreasing real valued sequence and $f=\inf\{f_{n}\in M(X,\mu):n\in\mathbb{N}\}$, then we write $f_{n}\downarrow f$. A Banach function space $L$ has an order continuous norm, if $0\leq f_{n}\downarrow 0$ implies $\|f_{n}\|_{L}\to 0$ as $n\to\infty$. It is well known that spaces $L^{p}(X,\mu)$, $1\leq p<\infty$, have order continuous norm. Moreover, the norm of any reflexive Banach function space is order continuous. In particular, we are interested in Banach function spaces $L$ such that $L$ and its Banach dual space $L^{*}$ have order continuous norms. Examples of such spaces are $L^{p}(X,\mu)$, $1<p<\infty$, while the space $L=c_{0}$ is an example of a non-reflexive Banach sequence space, such that $L$ and $L^{*}=l^{1}$ have order continuous norms.
By an operator on a Banach function space $L$ we always mean a linear operator on $L$. An operator $K$ on $L$ is said to be positive if it maps nonnegative functions to nonnegative ones, i.e., $KL_{+}\subset L_{+}$, where $L_{+}$ denotes the positive cone $L_{+}=\{f\in L:f\geq 0\;\mathrm{a.e.}\}$. Given operators $K$ and $H$ on $L$, we write $K\geq H$ if the operator $K-H$ is positive.
@@
where $D_{L}=\{f\in L:\|f\|_{L}\leq 1\}$. Then $\gamma(K)\leq\|K\|$, $\gamma(K+H)\leq\gamma(K)+\gamma(H)$, $\gamma(KH)\leq\gamma(K)\gamma(H)$ and $\gamma(\alpha K)=\alpha\gamma(K)$ for $\alpha\geq 0$. Also $0\leq K\leq H$ implies $\gamma(K)\leq\gamma(H)$ (see e.g. [23, Corollary 4.3.7 and Corollary 3.7.3]). Let $r_{ess}(K)$ denote the essential spectral radius of $K$, i.e., the spectral radius of the Calkin image of $K$ in the Calkin algebra. Then
@@
and $r_{ess}(K)\leq\gamma(K)$. Recall that if $L=L^{2}(X,\mu)$, then $\gamma(K^{*})=\gamma(K)$ and $r_{ess}(K^{*})=r_{ess}(K)$, where $K^{*}$ denotes the adjoint of $K$ (see e.g. [23, Proposition 4.3.3, Theorems 4.3.6 and 4.3.13 and Corollary 3.7.3], [27, Theorem 1], [22]). Note that equalities (2.2) and $r_{ess}(K^{*})=r_{ess}(K)$ are valid for any bounded operator $K$ on a given complex Banach space $L$ (see e.g. [23, Theorem 4.3.13 and Proposition 4.3.11], [27, Theorem 1]).
It is well-known that kernel operators play an important, often even central, role in a variety of applications from differential and integro-differential equations, problems from physics (in particular from thermodynamics), engineering, statistical and economic models, etc (see e.g. [18, 31] and the references cited there). For the theory of Banach function spaces and more general Banach lattices we refer the reader to the books [41, 4, 1, 2, 23].
Let $\Sigma$ be a bounded set of bounded operators on a complex Banach space $L$. For $m\geq 1$, let
$\Sigma^{m}=\{A_{1}A_{2}\cdots A_{m}:A_{i}\in\Sigma\}.$
The generalized spectral radius of $\Sigma$ is defined by
$r(\Sigma)=\limsup_{m\to\infty}\;[\sup_{A\in\Sigma^{m}}r(A)]^{1/m}$ (2.3)
and is equal to
$r(\Sigma)=\sup_{m\in\mathbb{N}}\;[\sup_{A\in\Sigma^{m}}r(A)]^{1/m}.$
The joint spectral radius of $\Sigma$ is defined by
$\hat{r}(\Sigma)=\lim_{m\to\infty}[\sup_{A\in\Sigma^{m}}\|A\|]^{1/m}.$ (2.4)
Similarly, the generalized essential spectral radius of $\Sigma$ is defined by
$r_{ess}(\Sigma)=\limsup_{m\to\infty}\;[\sup_{A\in\Sigma^{m}}r_{ess}(A)]^{1/m}$ (2.5)
and is equal to
$r_{ess}(\Sigma)=\sup_{m\in\mathbb{N}}\;[\sup_{A\in\Sigma^{m}}r_{ess}(A)]^{1/m}.$
The joint essential spectral radius of $\Sigma$ is defined by
$\hat{r}_{ess}(\Sigma)=\lim_{m\to\infty}[\sup_{A\in\Sigma^{m}}\gamma(A)]^{1/m}.$ (2.6)
It is well known that $r(\Sigma)=\hat{r}(\Sigma)$ for a precompact nonempty set $\Sigma$ of compact operators on $L$ (see e.g. [37, 38, 24]), in particular for a bounded set of complex $n\times n$ matrices (see e.g. [14, 36, 10, 26] and the references cited there). This equality is called the Berger-Wang formula or also the generalized spectral radius theorem. It is known that also the generalized Berger-Wang formula holds, i.e, that for any precompact nonempty set $\Sigma$ of bounded operators on $L$ we have
$\hat{r}(\Sigma)=\max\{r(\Sigma),\hat{r}_{ess}(\Sigma)\}$
(see e.g. [38, 24, 37]). Observe also that it was proved in [24] that in the definition of $\hat{r}_{ess}(\Sigma)$ one may replace the Hausdorff measure of noncompactness by several other seminorms, for instance it may be replaced by the essential norm.
In general $r(\Sigma)$ and $\hat{r}(\Sigma)$ may differ even in the case of a bounded set $\Sigma$ of compact positive operators on $L$ (see [36] or also [30]). Also, in [17] the reader can find an example of two positive non-compact weighted shifts $A$ and $B$ on $L=l^{2}$ such that $r(\{A,B\})=0<\hat{r}(\{A,B\})$. As already noted in [37] also $r_{ess}(\Sigma)$ and $\hat{r}_{ess}(\Sigma)$ may in general be different.
The theory of the generalized and the joint spectral radius has many important applications for instance to discrete and differential inclusions, wavelets, invariant subspace theory (see e.g. [10, 40, 37, 38] and the references cited there). In particular, $\hat{r}(\Sigma)$ plays a central role in determining stability in convergence properties of discrete and differential inclusions. In this theory the quantity $\log\hat{r}(\Sigma)$ is known as the maximal Lyapunov exponent (see e.g. [40]).
We have the following well known facts that hold for all $\rho\in\{r,\hat{r},r_{ess},\hat{r}_{ess}\}$:
$\rho(\Sigma^{m})=\rho(\Sigma)^{m}\;\;\mathrm{and}\;\;\rho(\Psi\Sigma)=\rho(\Sigma\Psi)$ (2.7)
where $\Psi\Sigma=\{AB:A\in\Psi,B\in\Sigma\}$ and $m\in\mathbb{N}$.
@@
Let us recall the following result, which was proved in [13, Theorem 2.2] and [28, Theorem 5.1 and Example 3.7] (see also e.g. [30, Theorem 2.1]).
@@
If, in addition, $L=L^{2}(X,\mu)$, then
$w(K)$$\leq$ $w\left((K_{11}\cdots K_{n1})^{(\alpha_{1})}\circ\cdots\circ(K_{1m}\cdots K_{nm})^{(\alpha_{m})}\right)$ (2.14)
$\leq$ $w(K_{11}\cdots K_{n1})^{\alpha_{1}}\cdots w(K_{1m}\cdots K_{nm})^{\alpha_{m}}.$
@@
If, in addition, $L=L^{2}(X,\mu)$, then
$w(K_{1}^{(\alpha_{1})}\circ K_{2}^{(\alpha_{2})}\circ\cdots\circ K_{m}^{(\alpha_{m})})\leq w(K_{1})^{\alpha_{1}}w(K_{2})^{\alpha_{2}}\cdots w(K_{m})^{\alpha_{m}}.$ (2.19)
Let $\Psi_{1},\ldots,\Psi_{m}$ be bounded sets of positive kernel operators on a Banach function space $L$ and let $\alpha_{1},\ldots\alpha_{m}$ be positive numbers such that $\sum_{i=1}^{m}\alpha_{i}=1$. Then the bounded set of positive kernel operators on $L$, defined by
$\Psi_{1}^{(\alpha_{1})}\circ\cdots\circ\Psi_{m}^{(\alpha_{m})}=\{A_{1}^{(\alpha_{1})}\circ\cdots\circ A_{m}^{(\alpha_{m})}:A_{1}\in\Psi_{1},\ldots,A_{m}\in\Psi_{m}\},$
is called the weighted Hadamard (Schur) geometric mean of sets $\Psi_{1},\ldots,\Psi_{m}$. The set $\Psi_{1}^{(\frac{1}{m})}\circ\cdots\circ\Psi_{m}^{(\frac{1}{m})}$ is called the Hadamard (Schur) geometric mean of sets $\Psi_{1},\ldots,\Psi_{m}$.
@@
$f_{1}^{\alpha}g_{1}^{\beta}+\cdots+f_{m}^{\alpha}g_{m}^{\beta}\leq(f_{1}+\cdots+f_{m})^{\alpha}(g_{1}+\cdots+g_{m})^{\beta}.$ (2.20)
@@
$(f_{11}^{\alpha_{1}}\cdots f_{1m}^{\alpha_{m}})+\cdots+(f_{n1}^{\alpha_{1}}\cdots f_{nm}^{\alpha_{m}})\leq(f_{11}+\cdots+f_{n1})^{\alpha_{1}}\cdots(f_{1m}+\cdots+f_{nm})^{\alpha_{m}}.$ (2.21)
The following result was established in [7, Theorems 3.2(i) and 3.6(i)] by applying Theorem 2.1 and (2.21).
Theorem 2.3.
Let $\{\Psi_{ij}\}_{i=1,j=1}^{k,m}$ be bounded sets of positive kernel operators on a Banach function space $L$, let $\rho\in\{r,\hat{r}\}$ and assume that $\alpha_{1},\ldots,\alpha_{m}$ are positive numbers such that $\sum_{i=1}^{m}\alpha_{i}=1$.
Then
$\rho\left(\left(\Psi_{11}^{(\alpha_{1})}\circ\cdots\circ\Psi_{1m}^{(\alpha_{m})}\right)\ldots\left(\Psi_{k1}^{(\alpha_{1})}\circ\cdots\circ\Psi_{km}^{(\alpha_{m})}\right)\right)$ (2.22)
$\leq$ $\rho\left((\Psi_{11}\cdots\Psi_{k1})^{(\alpha_{1})}\circ\cdots\circ(\Psi_{1m}\cdots\Psi_{km})^{(\alpha_{m})}\right)$
$\leq$ $\rho\left(\Psi_{11}\cdots\Psi_{k1}\right)^{\alpha_{1}}\cdots\rho\left(\Psi_{1m}\cdots\Psi_{km}\right)^{\alpha_{m}}$
and
$\rho\left(\left(\Psi_{11}^{(\alpha_{1})}\circ\cdots\circ\Psi_{1m}^{(\alpha_{m})}\right)+\ldots+\left(\Psi_{k1}^{(\alpha_{1})}\circ\cdots\circ\Psi_{km}^{(\alpha_{m})}\right)\right)$ (2.23)
$\leq$ $\rho\left((\Psi_{11}+\cdots+\Psi_{k1})^{(\alpha_{1})}\circ\cdots\circ(\Psi_{1m}+\cdots+\Psi_{km})^{(\alpha_{m})}\right)$
$\leq$ $\rho\left(\Psi_{11}+\cdots+\Psi_{k1}\right)^{\alpha_{1}}\cdots\rho\left(\Psi_{1m}+\cdots+\Psi_{km}\right)^{\alpha_{m}}.$
If, in addition, $L$ and $L^{*}$ have order continuous norms, then (2.22) and (2.23) hold also for $\rho\in\{r_{ess},\hat{r}_{ess}\}$.
@@
$\rho(S_{\alpha}(K_{1})+\ldots+S_{\alpha}(K_{n}))\leq\rho(K_{1}+\cdots+K_{n}).$ (2.25)
The following two results were proved in [8, Theorems 3.1 and 3.5].
@@
Let $K_{1},\ldots,K_{n}$ be positive kernel operators on $L=L^{2}(X,\mu)$. For $\rho\in\{r,r_{ess},\gamma,\|\cdot\|,w\}$ define $\rho_{n}:[0,1]\to[0,\infty)$ by